OpenStudy (anonymous):

Could someone walk me thru how to do this? (4m^2+8m)/4m^2

OpenStudy (emilylauren):

is this equation equal to anything?

OpenStudy (emilylauren):

\[(4m^{2}+8m)\div(4m^{2})\]

OpenStudy (anonymous):

No, not equation.

OpenStudy (emilylauren):

Are you just simplifying then? I would divide the fractions into two parts so that (4m^2/4m^2)+(8m/4m^2)

OpenStudy (emilylauren):

so then it simplifies to 1+(2/m)

OpenStudy (anonymous):

OH okay so 8/4 2, and then the m and the m^2 were subtracted out leaving only one m?

OpenStudy (emilylauren):

(4m^2/4m^2) is equal to the 1 (8m/4m^2) you can cancel out one m, so one is left on the bottom, and simplify the fraction 8/4 to 2

OpenStudy (emilylauren):

since there was an m in the numerator and the denominator I could just factor it out :)

OpenStudy (emilylauren):

m/m^2 = 1/m Does that make sense? You can just factor it out.

OpenStudy (anonymous):

Okay no. Lol. I get the other pieces for it but the variable part is still iffy.

OpenStudy (anonymous):

So if I am doing something like\[\frac{ 25a ^{5} }{ -5a }\] I would end up with \[-5a ^{4}\] ?

OpenStudy (anonymous):